the effective pH range of the KF/HF (Ka=6.8x10^-4) buffer is?
a) 2.2-4.2
b)1.2-4.2
c) 2.2-5.2
d) all possible
Answer #1

ka = 6.8x104 pka = 109 (6.8x109) - 4-0.83250 = 3.1674 pka : 3.167 = 3.17 pH range is pkat , pka_1 , 3.167+1 Pranze = 3.167-1